OpenStudy (anonymous):

Find the direction angle for v = <1, 4>. 14° 76° 104° 256°

OpenStudy (anonymous):

|v| = √(x² + y²) = √ ( (1)² + (4)² ) =

OpenStudy (anonymous):

arctan (1/ 4)=

OpenStudy (anonymous):

14.04

OpenStudy (anonymous):

and then I need to find the quadrant

OpenStudy (tkhunny):

 Does that look like 14º?

OpenStudy (anonymous):

so would it be 104?

OpenStudy (anonymous):

104º

OpenStudy (tkhunny):

 Try to pay better attention. Vertical is 90º. Is it past that? BTW \(atan(4) = \) what?